OpenStudy (anonymous):

Use substitution to solve for y in the system of equations. 4x-2y=56 5x+y=28 A.Y=3 B.Y=-12 C.Y=-18 D.Y=12

OpenStudy (ash2326):

@kaek98 Do you know substitution?

OpenStudy (anonymous):

I don't know much of it. I know you replace y with a number

OpenStudy (anonymous):

oh that's easy

OpenStudy (anonymous):

put y =28-5x in first equation solve for x then find y

OpenStudy (ash2326):

We 'll find x in terms of y from first equation, then substitute for x in the second. I'll illustrate that

OpenStudy (anonymous):

tell me the answer

OpenStudy (anonymous):

after solving

OpenStudy (ash2326):

\[4x-2y=56\] Add 2y to both the sides \[4x=56+2y\] Divide both sides by 4 \[x=\frac{56+2y}{4}\] Do you get this ? @kaek98

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

@ash2326 It does somewhat.

OpenStudy (ash2326):

Great now we have second equation \[5x+y=28\] Let's put the value of x found earlier \[5\times \frac{(56+2y)}{4}+y=28\] Multiply both sides by 4 \[4 \times (5\times \frac{(56+2y)}{4}+y)=4\times 28\] \[5\times (56+2y)+4y=112\] Can you try from here @kaek98

OpenStudy (anonymous):

I have this one. When I put in -12 I came up with 112. But I can't get the first one

OpenStudy (ash2326):

First one ?

OpenStudy (anonymous):

4x-2y=56

OpenStudy (ash2326):

We need to find y, you already found that y=-12 Now if you need to find x \[4x-2y=56\] put y=-12 here and solve for x

OpenStudy (anonymous):

x= 8

OpenStudy (ash2326):

yeah :) Good work :D

OpenStudy (anonymous):

@ash2326 Do you think you could help me with another one? I had someone helping me with it but they left

OpenStudy (ash2326):

yeah, post that as a new question :)

OpenStudy (anonymous):

Okay Thank you :)